OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL AND FAN the table shows the solution to the equation [2x - 4] - 3 = 3: step 1: 2x - 4| = -3 + 3 step 2: 2x - 4| = 0 step 3: 2x - 4 = 0 step 4: 2x = 4 step 5: x =2 which is the first incorrect step? A. step 1 B. step 2 C. step 3 D. solution is correct

OpenStudy (anonymous):

If you wrote the equation down correctly, then the first step is incorrect. |2x-4|-3=3 add 3 to both sides |2x-4|-3+3= 3+3 simplify |2x-4|=6